Project

General

Profile

NDOS Commissioning Meeting 2011/03/31

Attendance: Jon, Peter, Andrew, Gary, Mark Mathis, Mat, Leon, Susan, Jarek, Jan, Nick, Ralf, Minerba, Gavin, Denis, Athans

Agenda:

  1. NDOS Status
  2. DataCheck Studies
  3. DSO Studies
  4. DAQ Update & Plans
  5. DCS Status & Plans
  6. AOB

NDOS Status:

  • Mark Gephart's new nozzle head designed to top-off NDOS modules was found to not work quite as well
  • Mat:
    • Denis ran validateDCM scripts on 3 DCMs on di-block 3, and that went successfully
    • Today, FEBs were hung on top of 3rd di-block on all positions that had filled modules, as well as all modules on muon-catcher. Basically every other plane. Skipped first and last plane because there were problems with filling those. Denis will run the validateFEB on these.
  • Ralf: APDs? Leon: On their way.
  • DAQ stability: problems whenever anyone is doing work on detector or DAQ software.
  • Jon brought up a concern that people doing either hardware or software work are sometimes not leaving the "system" in a state that allows shifters to take data. Jon suggested that work should stop ~1 hour before the person doing the work wishes to leave for the day so they can "clean up after themselves" and ensure that we can get a run going. Not everyone liked the idea, in particular a hard 1 hour rule. Jon is reconsidering another approach to ensure that workers do proper clean-up...
  • Andrew: network problems are now understood and have been resolved. If problems arise again, please file a Help Desk ticket.

DataCheck Studies:

DSO Studies:

DAQ Update & Plans:

  • Andrew: we had a problem starting runs earlier today. Runs would begin and immediately crash. Finally traced it to the fact that the GPS unit has lost its lock to satellites but was reporting as fine. Techs believe that the reset is not resetting the GPS units. Possible to fix. In the end, we did a power cycle and it came back ok. We've been running fine since.
  • Peter:
    • missing trigger issue: prob. of missing a trigger is down to 1/subrun. Gennadiy report that this lost trigger is not due to DDS.
    • Sue has made some updates to Ganglia with new plotting features.
    • next week:
      • Ron will import a change to the DCMs that will eliminate network errors (Monday)
      • Rick K. would like to check timing ports on all DCMs. Will be disruptive to data taking. Test took some 10s of minutes at teststand. So this could be a couple of hours of work.
  • Mat: does it make sense to add DCMs with FEBs to timing chain on 3rd di-block? Peter's not sure if it makes sense, but we could in the least test that timing signals are being propagated. This should probably be added to the DCM validation procedure (requires new kernel driver)

DCS Status & Plans:

  • Chad:
    • RTDs are soldered and will be installed next week.
    • procured RJ45 plugs and will install at NDOS and patch panels
    • modified recording of DCS times to be UTC
    • power supplied are now being read out on both ends (sense and power)
    • local logging is also implemented
  • Athans:
    • able to read FEB temperatures on teststand
    • GUI is a work in progress

AOB